OpenStudy (anonymous):

Use a calculator to find the measure of acute angle theta to the nearest 10 minutes.

OpenStudy (anonymous):

\[\csc \theta=1.046\]

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

This is a similar situation like the last question. csc(theta) is the same as 1/sin(theta) Solve for sin(theta), then use arcsin function.

OpenStudy (anonymous):

I'm confused :/

OpenStudy (ipwnbunnies):

\[\csc(\theta) = \frac{1}{\sin \theta}\]

OpenStudy (anonymous):

am I right so far ...\[\sin \theta=0.9560\]

OpenStudy (ipwnbunnies):

\[\frac{1}{\sin \theta} = 1.046\]

OpenStudy (ipwnbunnies):

Correct, now use the arcsine function to figure out which angle has a sine of 0.9560.

OpenStudy (anonymous):

\[\theta=\sin^{-1} 0.9560=72.9404\]

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (anonymous):

\[\theta=73degrees60\min\]

OpenStudy (ipwnbunnies):

I'm not into minutes or seconds, sorry.

OpenStudy (anonymous):

oh okay,thanks for the help. I really don't know what the final answer is because I'm trying to figure out the minutes.

OpenStudy (ipwnbunnies):

Whoa, Idk what I just did. && Sorry man.